The longest side of a right triangle is 4 inches more than the shortest side. The third side is 8 inches. Find the length of each side.
What i've tried : x^2+8^2= (4x+x)^2 (Pythagorean theorem)
x^2+64= (4x+x) (4x+x)
x^2+64= 16x^2+4x^2+4x^2+x^2 (FOIL method)
x^2+64= 25x^2 (might've done the foiling wrong)
Then i subtracted x^2 and 64 to bring terms to the right and got a quadratic equation: 0=24x^2-64
Factored :(6x-8) (4x+8)
My solutions were 4/3 and 2. Not sure if this is correct... they don't match up when plugged into the pythagorean theorum. Any assist on this would be greatly appreciated. The examination will be presented on wed.

Answer by richwmiller(17219)

shortest side x
other side 8
longest side=hypotenuse x+4
x^2+64=(x+4)^2
x=6
x+4=10
6,8,10 the classic 3,4,5 right triangle